OpenStudy (anonymous):

How do I convert y=2x^2-4x+7 into vertex form?

OpenStudy (anonymous):

Vertex form: \[y = a(x - h)^{2} + k\]

OpenStudy (anonymous):

You have your equation in standard form.

OpenStudy (anonymous):

I think, it's in standard form right?

OpenStudy (anonymous):

I forgot that equation.

OpenStudy (anonymous):

Yes it is

OpenStudy (anonymous):

Cool, let me think...

OpenStudy (anonymous):

in standard form, x is the axis of symmetry, correct?

OpenStudy (anonymous):

Yeah, it is. hmmz

OpenStudy (anonymous):

\[x = -\frac{ b }{ 2a }\] is the axis of symmetry.

OpenStudy (anonymous):

\[y = 2x^{2} - 4x + 7\] \[x = -\frac{ b }{ 2a }\] \[x = -\frac{ -4 }{ 2(2) }\] \[x = -\frac{ -4}{ 4 }\]

OpenStudy (anonymous):

With me so far? Any questions?

OpenStudy (anonymous):

Possibly but hold on

OpenStudy (anonymous):

Whenever you're ready. :)

OpenStudy (anonymous):

Can you please check this

OpenStudy (anonymous):

One second, taking my dog outside.

OpenStudy (anonymous):

Okay, checking.

OpenStudy (anonymous):

Your work looks correct to me. :D

OpenStudy (anonymous):

Ok thank you

OpenStudy (anonymous):

No problem.

OpenStudy (the_fizicx99):

Y = 2x^2 - 4x + 7 I normally start by subtracting 7 from both sides, -7 = 2x^2 - 4x Now you can factor this: 2x^2 - 4x -> 2(x^2 - 2x) Then taking half the "x" term, and then squaring it. \(\ \sf -7 =2(x^2 - 2\color{blue}{x})\) So, \(\ \dfrac{-2}{2} = -1, -1^2 = 1 \) therefore, I add 1 into the parenthesis, -7 = 2([x^2 - 2x + 1] -1) ... you added 1 into the parenthesis, to keep it balanced you subtract 1 outside it. Now factor x^2 - 2 + 1, you'll always get (x \(\ \pm \) ? )^2 Where the "?" is, you'll always have half the "x" term, so -2/2 = -1, so: (x - 1)^2 is the factored form of x^2 - 2x + 1. -7 = 2[(x - 1)^2 - 1], distribute the 2 to -1, 2*-1 = -2 -7 = 2(x - 1)^2 - 2 Add 7 to the RHS, -7 = 2(x - 1)^2 - 2 +7 +7 y = 2(x - 1)^2 + 5 And that's vertex form, \(\ \Large y = a(x - h)^2 + k\) The vertex is at: (h,k) The axis of symmetry is the x coordinate in the vertex. Vertex: (1, 5) Axis: x = 1